October 23, 2014 FDC+CDC meeting
- Instructions for Bluejeans meeting connection
- FDC meeting ID: 290664653
- To join via a Web Browser, go to the page  https://bluejeans.com/290664653.
- DAQ (Beni)
- f125 status
- Electronics (Fernando)
- CDC update (Mike)
- FDC update (Vlad, Lubomir)
- Commissioning Wiki Page (all)
Participants: Curtis, Naomi (CMU) Mike, Fernando, Chris, Eugene, Tim, Dave, Beni, Nick, Simon, Vlad, Cody, Sergey, and Lubomir (JLab). .
DAQ and electronics
- Cody generated a new version of the firmware that is much more stable, as Beni tested it over the last night. With the new firmware Sasha was also able to have the CDC crates running at >1kHz without crushes for some time (before, they crushed at about 300Hz). However, with the new firmware, some new modules started showing problems, as tested by Beni and Sergey. A chip-scope was installed on the last module of ROCFDC5, but turns out other modules were crushing the DAQ now.
- Fernando explained the changes in the firmware: in the previous version, Cody let the CAD software do the planing of the changes (in particular going from 16 to 72 channels) to the f250 firmware. As a result, some of the clocks were not routed in an uniform way. Now Cody manually required a separate (clean) area to be used for the clock distribution.
- Cody: the new bad modules are most likely related to a different problem, could be a hardware problem. We are not confident in the quality of the boards (Fernando). The first article f125 from the new vendor will arrive next week.
- Eugene: For the first days of running we should simply exclude the bad modules from the DAQ till we find (if any) a stable configuration that runs for at least an hour without crushes. Then, during the several days of down time we may tray to fix/replace the bad modules. At least for the first days, the main priority will be to have the DAQ working in any configuration (even without the chambers!). Then we can think of how to test the detectors, which may be postponed till December.
- Bryan Moffit is working on a software reset of the module if it crushes, so that you re-sync the events and use the information from that module further without stopping the DAQ.
- Mike is working on the dac/pedestal calibration. Simple slope extrapolation doesn't work. He is using a standalone code with Beni's DAQ, that will be used for checking the rates when the main DAQ is not working.
- Tim: next week we will have the results from the last FDC survey, namely z-positions and plane angles, SEPARATELY for each package.
- Vlad showed plots with gain coefficients as reconstructed from the EEL125 data separately for each cathode (first plot linked above). Now he is simulating exclusion of a group of bad channels (like bad module) to see if the calibration of the rest is affected. Second plot shows the error in the coefficients for the different channels in case a group of channels (~ 160-170) is excluded.
- Simon did a plugin for the FDC online monitoring, showing the strip and wire profiles as discussed last time. Wire profiles were OK, but not the strips because of some noisy channels (due to wrong dac values for example). Lubomir continued working, adding more information to the code using position and timing correlation between the two cathodes and the wires to filter out the noise. While the position correlations are OK, there are inconsistencies with the timing correlations. Due to the lack of the trigger time (not yet implemented in JANA/DANA frame) Lubomir tried to synchronize the wire timing using the f125 timing information.